OpenStudy (anonymous):

evaluate the line integral, where C is the given curve:

OpenStudy (anonymous):

\[\int\limits_{}^{}y ^{3}dS\] C: x= t^3, y = t, 0<t<2

OpenStudy (anonymous):

i sep getting (1/36)(145^(3/2) - 1) and it should be (1/54)(145^(3/2) - 1)

OpenStudy (turingtest):

\[ds=\sqrt{(3t^2)^2+1}=\sqrt{9t^4+1}\]\[\int_{0}^{2}t^3\sqrt{9t^4+1}dt\]This should be the right setup I think

OpenStudy (turingtest):

@allyfranken yes, this gives the right answer let me know where you are having your trouble

OpenStudy (anonymous):

it gives you 1/54? bc i et 1/36 when i do it

OpenStudy (anonymous):

yeah i did the integral above and i got 1/36 instead of 1/54

OpenStudy (anonymous):

wouldn't ds be ||r'(t)|| not ||r(t)|| though?

OpenStudy (experimentx):

Looks like Turing was right after all.

OpenStudy (anonymous):

can you walk me through it bc i keep getting 1/36 instead of 1/54

OpenStudy (experimentx):

you meant he integral??

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

omg nvm i got it

OpenStudy (experimentx):

let t^2 = 1/3 tan x

OpenStudy (experimentx):

2t dt = 1/2 sec^2x dx

OpenStudy (experimentx):

dt = 1/4 sec^2x/sqrt(tanx)

OpenStudy (experimentx):

*edit* dt = 1/4 sec^2x sqrt(3)/sqrt(tanx) dx y^3 = tan^3/2x /3sqrt(3) y^3 dt = tan^(3/2)x /3sqrt(3) * 1/4 sec^2x sqrt(3)/sqrt(tanx) dx = tan^(3/2)x /3sqrt(3) * 1/4 sec^2x sqrt(3)/sqrt(tanx) dx = 1/12 * tanx sec^x dx

OpenStudy (experimentx):

what have you got??

OpenStudy (experimentx):

after putting that values in square root we get 1/12 * tanx sec^3x dx

OpenStudy (experimentx):

seems like I got the same http://www.wolframalpha.com/input/?i=integrate+1%2F12+tanx+sec%5E3x

OpenStudy (turingtest):

@experimentX why did you not just use\[u=9t^4+1\]? :P

OpenStudy (experimentx):

lol ... didn't see that.

OpenStudy (experimentx):

It could have made a lot simpler. since i've used that kind of method one or two times here on OS, i've not build insight. Well, a step to remember.

OpenStudy (turingtest):

just had to point it out simplicity=elegance